William opened two investment accounts for his grandson's college fund. The first year, these investments, which totaled $18,000 yielded $831 in simple interest. Part of the money was invested at 5.5% and the rest at 4%. How much ws invested at each rate?

Answer by checkley77(12844)

.055X+.04(18,000-X)=831
.055+720-.04X=831
.015X=831-720
.015X=111
X=111/.015
X=7,400 AMOUNT INVESTED @ 5.5%.
18,000-7,400=10,600 AMOUNT INVESTED @ 4%
PROOF:
.055*7,400+.04*10,600=831
407+424=831
831=831